OpenStudy (anonymous):

The equation |4x + 2| + 3 = 2 has no solution. o.O THIS TOTALLY FALSE RIGHT

OpenStudy (bahrom7893):

yes

OpenStudy (anonymous):

it has not solution because the absolute value cannot be negative

OpenStudy (hoblos):

this is true

Directrix (directrix):

Yes, it is 100% wrong.

OpenStudy (bahrom7893):

wait i meant false as in the statement is not true

Directrix (directrix):

The equation |4x + 2| + 3 = 2 has no solution. TRUE statement

OpenStudy (bahrom7893):

your statement china.. lol direct is right

OpenStudy (anonymous):

its a TRUE OR FALSE QUESTION LOL I'M SAYIN IT'S FALSE

OpenStudy (anonymous):

but waittttt o.O so its true?

OpenStudy (hoblos):

<<The equation |4x + 2| + 3 = 2 has no solution>> this is true

Directrix (directrix):

The equation |4x + 2| + 3 = 2 has no solution. TRUE statement

Directrix (directrix):

><)))*>

OpenStudy (anonymous):

lol!! ok